OpenStudy (anonymous):

Which rational number equals 0 point 1 with bar over 1? 1 over 11 1 over 10 1 over 9 1 over 8

OpenStudy (welshfella):

10 * 0.111... 1 =1 1111... 1 ** 0.111.. = 0 . 11111. subtract 9 * 0.11.. = 1 so 0.111... = 1/9